Project

General

Profile

10/22/2013 Minutes

Attendance

Michael (has peanut M&Ms), Nick, Mat, Gavin, Gareth, Sergey, Athans, Kanika, Dominick, Susan, Evan

We decided to hold daily meetings at 1 pm-2 pm. We will try to book a room.

Wiki-page - Shifter (good)
- Shifter run list
- Issues


Operations

Susan reported NDOS is running relatively fine.
FD is running cold and with non mangled cooling.

Production Runs

Data taking -> Raw2Root 24 hrs
Script to pick up new runs (every 3 hours) (24hrs)
Fast tracking the runs -> We will monitor the situation

Issues

-Early data issue type

Dom - ~90% cosmic trigger caf jobs crash
- W methods are dividing by small or zeros
Gav - newer tag - data doesn't store the gdml data
- Talk to the sim group
- What is the default gdml? (The current tag picks up staggered geometry)
- Module stiffing.
DAQ - Trigger name changes. New configurations.
- Also the scripts parse the names now.
Nick - Mangled data issues.
- Hit efficiency (Warm fardet, G100), looked at hit efficiency.
- Fluor in scintillator, attenuation model
- Cold Far detector data. (Fraction of cold detector metric)
Dom - Cosmic rate in Data in much higher in data
- 11342 - Kanika
- Time duration of slices
- Get Run info (subset)


Kanika has agreed to be the first daily analysis "shifter".